# Disturbance-Adaptive Data-Driven Predictive Control: Trading Comfort Violation for Savings in Building Climate Control

### v3

@@ IV-B2 Setup of DAD-DPC
For the output $y$, Case 3 includes five dimensions corresponding to the five zones. The external input $w$ comprises the ambient dry-bulb temperature at ground level and global horizontal irradiation, as provided by the BOPTEST forecast and results APIs. For the control input $u$, heating power was selected for Cases 2 and 3, as the building thermodynamics can be reasonably approximated using linear models, making them suitable for RB-DPC. In Case 1, although nonlinearity arises from the HP’s coefficient of performance, RB-DPC effectively approximates short-term behaviors and benefits from adaptive updates for long-term operations [shi2025adaptive].
@@
Across all three cases, the following common parameters were used for DAD-DPC: sampling time, 15 minutes; prediction horizon $N$, 96 steps; initial steps $t_{init}$, 12 steps; regularization weight $Q_{g}$, 0.01II; regularization weight $Q_{\delta}$, 10; Hankel matrix data length $T$, 672 steps (one week); SCP data length $T_{c}$, 672 steps (one week); $\mathcal{Y}_{\text{lim}},\{y|{15}^{\circ}C\leq y\}$; $\alpha_{0}$, 0; update rate $\eta$, 0.5. The sampling time and the prediction horizon were selected based on our previous control experience in the BOPTEST challenge. They are also consistent with the practical MPC guideline for buildings from [drgovna2020all]: it recommends 10102020 samples over the rise time and reports typical ranges of 1515180180 minutes in the literature; the prediction horizon $N$ should be chosen so that $N$ sampling times span the system’s settling time while balancing model and forecast uncertainty. The parameters, $t_{init},Q_{g},T$ were chosen based on empirical guidelines from [shi2025adaptive]. A segmented trajectory trick from [o2022data] was applied in DPC to reduce data requirements. Two weeks of data for $T$ and $T_{c}$ were collected using a bang-bang controller for each case. A small $\alpha_{0}$ was selected to limit frequent violations at the beginning, while a moderate update rate $\eta$ was chosen to prevent extreme adaptive adjustments. Apart from those choices, in the RB-DPC, the control objective was set to $I(y_{t},u_{t})=u_{t}$, with input constraints $\mathcal{U}$ defined as the most recent electrical or heating power levels when the respective HVAC system was on. Under these settings, the resulting RB-DPC controllers were QP problems, as analyzed in Remark 2, and were all solved by Gurobi [gurobi] in all simulation cases. In this simulation study, various $\alpha$ values were tested to validate the satisfaction of the violation bound (2) using DAD-DPC. The results of this validation are presented in Section IV-C.
@@ IV-C Results of simulation cases
To evaluate the trade-off between violations and energy consumption, we tested the DAD-DPC with more $\alpha$ values and compared its performance with the default controller and the nominal DPC (i.e., $\pi(z_{t},\mathcal{D}(1))$, corresponding to $d_{i}\in\mathcal{D}_{i}(\sigma)=\{\mathbf{0}\}$ in the RB-DPC (8)). For all controllers except the default, five Monte Carlo simulations were conducted using different realizations of measurement noise. Figures 7, 8 and 9 display scatter plots of total relative energy consumption versus average violation level. Each marker shape represents a specific controller, and each data point corresponds to one Monte Carlo trial. For the DAD-DPC controller, each color indicates a different choice of $\alpha$. Table V summarizes the Monte Carlo averages of total relative energy consumption, average violation levels, reported in both ratio and magnitude, for each $\alpha$. Violation magnitude is computed as the product of the temperature deviation from the comfort bound and the sampling time, expressed in Kelvin-hours [Kh]. This metric serves as an additional comfort violation criterion, as suggested by the European indoor environment standard [comite2007indoor].
@@
The above results clearly illustrate the trade-off between comfort violations (both in ratio and magnitude) and energy consumption: lower $\alpha$ values reduced average violations and increased energy consumption. Compared to the conservative default controller for Cases 1, 2 and 3, DAD-DPC with a 5% average violation level, aligned with the European standard EN15251 Annex G [comite2007indoor], reduced energy consumption by 30.1%/11.2%/27.1%, respectively, demonstrating significant efficiency improvements. Furthermore, when compared to the nominal DPC, which let to more violations, the 5%-violation DAD-DPC setting resulted in only a marginal increase in energy consumption (2.8%/3.7%/3.7%) while achieving significantly fewer violations (77.6%/79.6%/73.2%) across the three cases. These findings highlight the ability of DAD-DPC to balance comfort and energy efficiency effectively.
While both violation ratio and magnitude showed trade-off behavior with energy consumption, they showed different patterns across the three cases. In all cases, the violation ratio approached the specified bound $\alpha$, but the violation showed clear variation. For example, when $\alpha=5\%$, the violation ratios were consistently close to the target, but the violation magnitudes were 0.030.03Kh, 0.880.88Kh, and 0.650.65Kh for the three cases, respectively. This is because, in the current DAD-DPC framework, the adaptive update of $\alpha_{t}$ in (LABEL:eqn:dad_alpha) is based on the binary violation indicator $v_{t}$, leading to feedback on violation ratio, as discussed in Section III-C. The authors of [korda2014stochastic] achieved different types of violation bounds by redefining $v_{t}$, assuming the disturbance distribution is known. Extending the DAD-DPC framework to account for violation magnitude using a similar approach is a promising direction for future research.
@@ V-A Setup in an occupied campus building
Computer: A central computer (iMac with 32 GB 1600 MHz DDR3 memory and a 3.4 GHz Intel Core i7) manages several components. (1) Database: Data, including HP power consumption, indoor temperature, and historical weather data, are stored in a time-series database, InfluxDB 1.3.7 [influxdb]. (2) DAD-DPC: The DAD-DPC framework was implemented in MATLAB. It calculated the control variables and communicated them to the HP using the Modbus serial communication protocol [swales1999open]. (3) Prediction Module: Weather forecasts, including outdoor air temperature and solar radiation, were obtained via the Tomorrow.io [tomorrow] API. Forecast data and current weather conditions were updated every 15 minutes.
Sensors: The HP’s power consumption was monitored using an EMU 3-phase meter [emu], which transmitted data to the central computer via Ethernet. Indoor temperature was monitored using four Fibaro Door/Window Sensor v2 units, installed at a height of 1.5 meters on different walls (locations shown in Figure 11). These sensors transmitted temperature data to the central computer every five minutes using the Z-wave wireless protocol. For this study, the building’s output, $y$, was calculated as the average temperature measured by the four sensors.
@@ V-B2 Setup of DAD-DPC
The parameters for the DAD-DPC method were chosen as follows: sampling time, 15 minutes; prediction horizon $N$, 24 steps; initial steps $t_{init}$, 12 steps; regularization weight $Q_{g}$, 0.01II; regularization weight $Q_{\delta}$, 10. The sampling time and prediction horizon were selected based on our previous control experience in the Polydome. They are also consistent with the practical MPC guidelines summarized in Section IV.B.2). The choice of $N=24$ (6 hours) was further constrained by the weather forecast API, Tomorrow.io [tomorrow], provided up to a 6-hour prediction in the 15-min sampling time. The segmented trajectory trick from [o2022data] was employed to reduce the data requirement for Hankel matrices. In the RB-DPC, the objective was defined as $I(y_{t},u_{t})=u_{t}$ and the input constraint $\mathcal{U}$ was set to the latest maximal electrical power. Under these settings, the resulting RB-DPC was a QP problem as analyzed in Remark 2, and was solved by MATLAB Quadprog [matlab:quadprog].
The experimental study spanned over two months and included various operational settings. The comfort constraints and parameters ($\alpha,\alpha_{0}$, $\eta$) were adjusted as described in Section V-C. To address the nonlinearity in the HP’s coefficient of performance, adaptive updates were incorporated into the RB-DPC [shi2025adaptive], and residual updates were applied to SCP for long-term operations. Initially, both the Hankel matrix and SCP data lengths were set to 480 steps (5 days). The required 10-day datasets for each mode were collected using the default controller.
@@ V-C Results of experiments
Average Violation
Ratio [%]
Average Violation
Magnitude [Kh]
Average Hourly Energy
Consumption [kWh]
Average Outdoor
Temperature [C]
Average Solar
Radiation [kW//m2]

### v4

@@ IV-B2 Setup of DAD-DPC
For the output $y$, Case 3 includes five dimensions corresponding to the five zones. The external input $w$ comprises the ambient dry-bulb temperature at ground level and global horizontal irradiation, as provided by the BOPTEST forecast and results APIs. For the control input $u$, heating power was selected for Cases 2 and 3, as the building thermodynamics can be reasonably approximated using linear models, making them suitable for RB-DPC. In Case 1, although nonlinearity arises from the HP’s coefficient of performance, RB-DPC effectively approximates short-term behaviors and benefits from adaptive updates for long-term operations [43].
@@
Across all three cases, the following common parameters were used for DAD-DPC: sampling time, 15 minutes; prediction horizon $N$, 96 steps; initial steps $t_{init}$, 12 steps; regularization weight $Q_{g}$, 0.01II; regularization weight $Q_{\delta}$, 10; Hankel matrix data length $T$, 672 steps (one week); SCP data length $T_{c}$, 672 steps (one week); $\mathcal{Y}_{\text{lim}},\{y|{15}^{\circ}C\leq y\}$; $\alpha_{0}$, 0; update rate $\eta$, 0.5. The sampling time and the prediction horizon were selected based on our previous control experience in the BOPTEST challenge. They are also consistent with the practical MPC guideline for buildings from [15]: it recommends 10102020 samples over the rise time and reports typical ranges of 1515180180 minutes in the literature; the prediction horizon $N$ should be chosen so that $N$ sampling times span the system’s settling time while balancing model and forecast uncertainty. The parameters, $t_{init},Q_{g},T$ were chosen based on empirical guidelines from [43]. A segmented trajectory trick from [34] was applied in DPC to reduce data requirements. Two weeks of data for $T$ and $T_{c}$ were collected using a bang-bang controller for each case. A small $\alpha_{0}$ was selected to limit frequent violations at the beginning, while a moderate update rate $\eta$ was chosen to prevent extreme adaptive adjustments. Apart from those choices, in the RB-DPC, the control objective was set to $I(y_{t},u_{t})=u_{t}$, with input constraints $\mathcal{U}$ defined as the most recent electrical or heating power levels when the respective HVAC system was on. Under these settings, the resulting RB-DPC controllers were QP problems, as analyzed in Remark 2, and were all solved by Gurobi [21] in all simulation cases. In this simulation study, various $\alpha$ values were tested to validate the satisfaction of the violation bound (2) using DAD-DPC. The results of this validation are presented in Section IV-C.
@@ IV-C Results of simulation cases
To evaluate the trade-off between violations and energy consumption, we tested the DAD-DPC with more $\alpha$ values and compared its performance with the default controller and the nominal DPC (i.e., $\pi(z_{t},\mathcal{D}(1))$, corresponding to $d_{i}\in\mathcal{D}_{i}(\sigma)=\{\mathbf{0}\}$ in the RB-DPC (8)). For all controllers except the default, five Monte Carlo simulations were conducted using different realizations of measurement noise. Figures 7, 8 and 9 display scatter plots of total relative energy consumption versus average violation level. Each marker shape represents a specific controller, and each data point corresponds to one Monte Carlo trial. For the DAD-DPC controller, each color indicates a different choice of $\alpha$. Table V summarizes the Monte Carlo averages of total relative energy consumption, average violation levels, reported in both ratio and magnitude, for each $\alpha$. Violation magnitude is computed as the product of the temperature deviation from the comfort bound and the sampling time, expressed in Kelvin-hours [Kh]. This metric serves as an additional comfort violation criterion, as suggested by the European indoor environment standard [11].
@@
The above results clearly illustrate the trade-off between comfort violations (both in ratio and magnitude) and energy consumption: lower $\alpha$ values reduced average violations and increased energy consumption. Compared to the conservative default controller for Cases 1, 2 and 3, DAD-DPC with a 5% average violation level, aligned with the European standard EN15251 Annex G [11], reduced energy consumption by 30.1%/11.2%/27.1%, respectively, demonstrating significant efficiency improvements. Furthermore, when compared to the nominal DPC, which let to more violations, the 5%-violation DAD-DPC setting resulted in only a marginal increase in energy consumption (2.8%/3.7%/3.7%) while achieving significantly fewer violations (77.6%/79.6%/73.2%) across the three cases. These findings highlight the ability of DAD-DPC to balance comfort and energy efficiency effectively.
While both violation ratio and magnitude showed trade-off behavior with energy consumption, they showed different patterns across the three cases. In all cases, the violation ratio approached the specified bound $\alpha$, but the violation showed clear variation. For example, when $\alpha=5\%$, the violation ratios were consistently close to the target, but the violation magnitudes were 0.030.03Kh, 0.880.88Kh, and 0.650.65Kh for the three cases, respectively. This is because, in the current DAD-DPC framework, the adaptive update of $\alpha_{t}$ in (LABEL:eqn:dad_alpha) is based on the binary violation indicator $v_{t}$, leading to feedback on violation ratio, as discussed in Section III-C. The authors of [27] achieved different types of violation bounds by redefining $v_{t}$, assuming the disturbance distribution is known. Extending the DAD-DPC framework to account for violation magnitude using a similar approach is a promising direction for future research.
@@ V-A Setup in an occupied campus building
Computer: A central computer (iMac with 32 GB 1600 MHz DDR3 memory and a 3.4 GHz Intel Core i7) manages several components. (1) Database: Data, including HP power consumption, indoor temperature, and historical weather data, are stored in a time-series database, InfluxDB 1.3.7 [23]. (2) DAD-DPC: The DAD-DPC framework was implemented in MATLAB. It calculated the control variables and communicated them to the HP using the Modbus serial communication protocol [46]. (3) Prediction Module: Weather forecasts, including outdoor air temperature and solar radiation, were obtained via the Tomorrow.io [47] API. Forecast data and current weather conditions were updated every 15 minutes.
Sensors: The HP’s power consumption was monitored using an EMU 3-phase meter [17], which transmitted data to the central computer via Ethernet. Indoor temperature was monitored using four Fibaro Door/Window Sensor v2 units, installed at a height of 1.5 meters on different walls (locations shown in Figure 11). These sensors transmitted temperature data to the central computer every five minutes using the Z-wave wireless protocol. For this study, the building’s output, $y$, was calculated as the average temperature measured by the four sensors.
@@ V-B2 Setup of DAD-DPC
The parameters for the DAD-DPC method were chosen as follows: sampling time, 15 minutes; prediction horizon $N$, 24 steps; initial steps $t_{init}$, 12 steps; regularization weight $Q_{g}$, 0.01II; regularization weight $Q_{\delta}$, 10. The sampling time and prediction horizon were selected based on our previous control experience in the Polydome. They are also consistent with the practical MPC guidelines summarized in Section IV.B.2). The choice of $N=24$ (6 hours) was further constrained by the weather forecast API, Tomorrow.io [47], provided up to a 6-hour prediction in the 15-min sampling time. The segmented trajectory trick from [34] was employed to reduce the data requirement for Hankel matrices. In the RB-DPC, the objective was defined as $I(y_{t},u_{t})=u_{t}$ and the input constraint $\mathcal{U}$ was set to the latest maximal electrical power. Under these settings, the resulting RB-DPC was a QP problem as analyzed in Remark 2, and was solved by MATLAB Quadprog [32].
The experimental study spanned over two months and included various operational settings. The comfort constraints and parameters ($\alpha,\alpha_{0}$, $\eta$) were adjusted as described in Section V-C. To address the nonlinearity in the HP’s coefficient of performance, adaptive updates were incorporated into the RB-DPC [43], and residual updates were applied to SCP for long-term operations. Initially, both the Hankel matrix and SCP data lengths were set to 480 steps (5 days). The required 10-day datasets for each mode were collected using the default controller.
@@ V-C Results of experiments
Average Violation Ratio [%] Average Violation Magnitude [Kh] Average Hourly Energy Consumption [kWh] Average Outdoor Temperature [C] Average Solar Radiation [kW//m2]